Suppose an investment offers to quintuple your money in 42 months (dont believe it). What rate of return per quarter are you being offered? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Rate of return 12.18-/-1%1%
Answer #1

We use the formula:
A=P(1+r/400)^4n
where
A=future value($5x )
P=present value($x say)
r=rate of interest
n=time period.(42/12=3.5 years)

5x=x*(1+r/400)^(4*3.5)

5=(1+r/400)^14

(5)^(1/14)=(1+r/400)

(1+r/400)=1.1218

r=1.1218-1

=12.18%(Approx).